# CVT-Occ: Cost Volume Temporal Fusion for 3D Occupancy Prediction

### v2

@@ 4.3 Main Results
As shown in Figure 1, We categorize previous temporal fusion methods into three types: (a) Attention-based methods, which adopt temporal self-attention to recurrently fuse the history BEV information; (b) WrapConcat-based methods, involving the alignment of BEV feature maps across different time steps; and (c) Plane-Sweep-based methods, which construct a cost volume for image pairs to utilize geometric constraints. Additionally, we introduce a novel approach, CVT-Occ, that utilizes geometric correspondences of 3D voxels over time to improve occupancy prediction accuracy. For comparative analysis, we extended some mainstream BEV models for the 3D semantic occupancy prediction task. We choose BEVFormer [20], BEVFormer-WarpConcat [37], and SOLOFusion [25] as representative methods for categories (a), (b), and (c), respectively.
For both BEVFormer [20] and SOLOFusion [25], we have replaced their original transformer decoders for object detection with a 3D occupancy prediction decoder. Additionally, for fairness, we omit the depth supervision module when reproducing SOLOFusion [25]. For BEVFormer-WrapConcat, we adopt a commonly used fusion method known as the Wrap and Concatenate strategy, as seen in BEVFormerv2 [37] and PanoOcc [35]. This strategy involves 3D coordinate alignment in BEV space based on the transformation matrix between historical and current frames. Subsequently, it concatenates previous BEV features with current BEV features along the channel dimension and applies residual blocks for temporal fusion. We incorporate the residual block proposed by PanoOcc [35] to create the BEVFormer-WrapConcat method. In the “BEVFormer-w/o TSA” configuration, we maintain the original architecture of BEVFormer [20] but exclude the Temporal Self-Attention(TSA) layers, thereby eliminating temporal attention mechanisms. In Table 1, we present the evaluation results for 3D occupancy prediction on the Occ3D-Waymo validation set. To ensure a fair comparison, all methods were implemented using the same network architecture and training strategy.

### v3

@@ 4.3 Main Results
As shown in Figure 1, We categorize previous temporal fusion methods into three types: (a) Attention-based methods, which adopt temporal self-attention to recurrently fuse the history BEV information; (b) WarpConcat-based methods, involving the alignment of BEV feature maps across different time steps; and (c) Plane-Sweep-based methods, which construct a cost volume for image pairs to utilize geometric constraints. Additionally, we introduce a novel approach, CVT-Occ, that utilizes geometric correspondences of 3D voxels over time to improve occupancy prediction accuracy. For comparative analysis, we extended some mainstream BEV models for the 3D semantic occupancy prediction task. We choose BEVFormer [20], BEVFormer-WarpConcat [37], and SOLOFusion [25] as representative methods for categories (a), (b), and (c), respectively.
For both BEVFormer [20] and SOLOFusion [25], we have replaced their original transformer decoders for object detection with a 3D occupancy prediction decoder. Additionally, for fairness, we omit the depth supervision module when reproducing SOLOFusion [25]. For BEVFormer-WarpConcat, we adopt a commonly used fusion method known as the Warp and Concatenate strategy, as seen in BEVFormerv2 [37] and PanoOcc [35]. This strategy involves 3D coordinate alignment in BEV space based on the transformation matrix between historical and current frames. Subsequently, it concatenates previous BEV features with current BEV features along the channel dimension and applies residual blocks for temporal fusion. We incorporate the residual block proposed by PanoOcc [35] to create the BEVFormer-WarpConcat method. In the “BEVFormer-w/o TSA” configuration, we maintain the original architecture of BEVFormer [20] but exclude the Temporal Self-Attention(TSA) layers, thereby eliminating temporal attention mechanisms. In Table 1, we present the evaluation results for 3D occupancy prediction on the Occ3D-Waymo validation set. To ensure a fair comparison, all methods were implemented using the same network architecture and training strategy.